Timing distribution
Slow Controls infrastructure and GUI
Vertical Slice preparedness
Online algorithm benchmark

Priorities with Lab Access:

Characterize Timing Stability (10-DTC-daisy chain or star-fanout-topology)
Debug DTC Event Building
Horizontal Slice Testing (detector subsystem priority is Vertical Slice Testing)
Retransmission testing


1. General status updates/Schedule review.

2. Status of DTC to ROC interface.

3. Status of EPICS on DAQ cluster.

4. Status of core ROC and timing distribution testing.

Rick -- not in attendance, has been working with Iris and Jason on 10 DTCs timing distribution

Eric -- new artdaq 3.9.0 release. Then ots and mu2e releases. Can build all software on CentOS8 on mu2edaq16. Need to pcie_ to see if any issues.

Ron -- Resurrected disk write rate tests. Working on a presentation (should be ready for next meeting). Preliminary results ~250MBps. Should try daq15.

Antonio -- working on slow controls widgets and opening pages (link navigation). Involved in snowmass 21 Mu2eII prep.

Bertrand -- In a week or two, will get back to optimization with Dave Brown.
Vivian -- new hardware for FCC test stand. Will move daq9 to lab 3 for VST in mid-August. Could try move with Material Move or through Jason/PREP cart/van. Thinking about structure/documentation of DCS - will propose something in EPICS meeting. Evan successfully had DQM modules working generating histograms. VST for Tracker mid-September.

Ryan -- new hire starting next week to help with distractions. We are getting new web cam at test stand.

Iris -- test stand ots configuration to setup 1 CFO + 10 DTCS across 6 nodes. Able to configure 1 + 8 DTCS (2 were down). After 4th or 5th link clock degradation was apparent. 8th was not staying locked. Gary would be interested in plots.